OpenStudy (anonymous):

taking the square root and show steps 4x^2+8x+3=0

OpenStudy (freckles):

sounds like you want to complete the square on one side to solve

OpenStudy (freckles):

\[4x^2+8x=-3 \\ \text{ divide both sides by 4 } \\ x^2+2x=-\frac{3}{4}\] do you recall how to factor x^2+2x+1?

OpenStudy (freckles):

hint: \[x^2+2x+1=(x+1)^2 \]

OpenStudy (freckles):

add one on both sides and try to use that last equation I posted

OpenStudy (anonymous):

4x^2 + 8x + 4 = 1 (2x+2)^2 = 1 2x+2 = plus or minus 1 x = -1/2 or -3/2 ?

OpenStudy (freckles):

seems great! \[x^2+2x=\frac{-3}{4} \\ x^2+2x+1=\frac{-3}{4}+1 \\ x^2+2x+1=\frac{-3}{4}+\frac{4}{4} \\ (x+1)^2=\frac{1}{4} \\ x+1= \pm \frac{1}{2} \\ x=- 1 \pm \frac{1}{2} =\frac{-2}{2} \pm \frac{1}{2}=\frac{-2 \pm 1}{2} \\ \text{ so } x=\frac{-2+1}{2}=\frac{-1}{2} \text{ or } x=\frac{-2-1}{2}=\frac{-3}{2}\] your way looks cute too :)

OpenStudy (anonymous):

thank you could you help me with completing the square on the same question?

OpenStudy (freckles):

on the same question you already did that

OpenStudy (freckles):

you wrote 4x^2+8x+4 as (2x+2)^2 (2x+2)^2 is the square of (2x+2)

OpenStudy (anonymous):

so taking the square and compting the square is the sam eequation?

OpenStudy (freckles):

I think I'm not sure what you are asking? you solved the equation above by completing the square.. then you took square root of both sides

OpenStudy (anonymous):

so im suppose to show my work for completing the square and taking the root but in that i did both right?

OpenStudy (freckles):

\[(2x+2)^2 =1 \text{ this is what you got after completing the square } \\ \text{ then you took square root of both sides } \\ \sqrt{(2x+2)^2}=\sqrt{1} \\ \sqrt{(2x+2)^2}=1 \ \text{ but this means } 2x+2=\pm 1 \]

OpenStudy (anonymous):

so thats the answer for completing the square ?

OpenStudy (freckles):

that is the answer by using the completing the square method

OpenStudy (freckles):

that is the answer you would also get by using any method there is to solve a quadratic such as the quadratic formula or factoring

OpenStudy (freckles):

though it would be a bit tricky to factor the expression on the left hand side since it wouldn't be doable over integers but that doesn't mean it can't be done at all

OpenStudy (anonymous):

could i give that whole equation and answer all the questions in one

OpenStudy (anonymous):

welll besides the factoring

OpenStudy (freckles):

I don't see how taking the square root and completing the square should have there on separate method you complete the square then take square root but anyways quadratic formula is easy you just use: \[x=\frac{-b \pm \sqrt{b^2-4ac}}{2a}\] factoring 4x^2+8x+3 is actually not tricky (I lied ) 4(3)=6(2) and 6+2=8

OpenStudy (freckles):

\[4x^2+8x+3 = 4x^2+6x+2x+3 \text{ factor this by grouping }\]

OpenStudy (anonymous):

ur confusing me with all this

OpenStudy (freckles):

but in either method you use you should get the solutions: x=-1/2 or x=-3/2

OpenStudy (freckles):

with the use of quadratic formula or factoring?

OpenStudy (anonymous):

both. ur talking about factoring rn or

OpenStudy (freckles):

what is rn?

OpenStudy (anonymous):

right now

OpenStudy (freckles):

I was talking about both quadratic formula and factoring above:

OpenStudy (freckles):

\[ax^2+bx+c=0 \text{ gives } x=\frac{-b \pm \sqrt{b^2-4ac}}{2a} \text{ is quadratic formula }\] you never seen this formula before?

OpenStudy (anonymous):

no i have but im not sure how it works

OpenStudy (freckles):

can you compare: ax^2+bx+c=0 to 4x^2+8x+3=0 and identify a,b, and c?

OpenStudy (anonymous):

my math program is super confusing and dosnt explain why it is

OpenStudy (anonymous):

no i just need to answer these question

OpenStudy (freckles):